# Evaluating Automatic Speech Recognition Systems for Korean Meteorological Experts

### v2

@@ 6 Conclusion
This work constructs a specialized evaluation dataset for assessing ASR models in the Korean meteorological domain, addressing the lack of domain-specific benchmarks. Our results show that pre-trained ASR models struggle with meteorological terminology, and while fine-tuning on general-domain data improves accuracy, domain adaptation remains essential for handling specialized vocabulary. While our dataset provides a valuable benchmark, future work should expand it to cover more diverse speech conditions, such as spontaneous and noisy environments, to better reflect real-world forecasting scenarios. By making this dataset available, we aim to drive further advancements in domain-adapted ASR systems for meteorological applications.

### v3

@@ 6 Conclusion
We introduce a specialized evaluation dataset for assessing ASR performance in the Korean meteorological domain, addressing the lack of domain-specific benchmarks. Our experiments reveal that while fine-tuning on general-domain data improves overall accuracy, specialized terminology remains a major source of error. By releasing this dataset, we aim to support further research in developing robust, domain-adapted systems that better reflect the demands of real-world forecasting scenarios.
Limitations
While our study provides a targeted benchmark for ASR in the Korean meteorological domain, several limitations remain. First, the dataset consists of scripted queries recorded in clean environments, which may not fully represent the acoustic variability found in real-world settings. In practice, meteorologists often speak spontaneously, with disfluencies, hesitations, or overlapping speech, especially during live broadcasts or team discussions. Our dataset does not yet capture such spontaneous or noisy conditions. Second, while we explored text-only post-processing using unimodal LLMs, we did not investigate more advanced correction strategies with acoustic cues or multiple ASR hypotheses. These approaches could further improve recognition of subtle or ambiguous terms but require additional data and computational resources. Lastly, our work is focused on a single domain (meteorology) and a single language (Korean). The findings may not generalize to other specialized domains, such as healthcare or law, or to other low-resource languages. Expanding this research to cross-domain and cross-lingual settings remains an important direction for future work.
Ethical Statement
All spoken queries in our dataset were recorded with the informed consent of native Korean speakers, who were compensated fairly for their participation. The recordings were collected in controlled environments to ensure quality and to minimize any unintended background content. To protect speaker privacy, no personally identifiable information (PII) was included in the dataset. All utterances were manually reviewed to avoid harmful, offensive, or culturally insensitive content. While the dataset aims to represent meteorological language use in Korean, we acknowledge that it may not fully capture regional or situational variations in speech patterns.
Acknowledgement
This work is supported by a grant of a Developing Intelligent Assistant Technology and Its Application for Weather Forecasting Process (KMA2021-00123), Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)), and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2025-00555621).